mod use Show that the following equations have no integer solutions: (4) 25 + 12x2 + 24x + 1 = 0.
Answer #1

25 + 122?+ 24x+1 = 0 25 + 122? +243 = -1 (mod 12) -15ll(mod 12) 3 equations des ...x = 11 (modie) Thus, the +123*+24x+1 -0 do